# Persistent Backdoor Attacks in Continual Learning

### v1

@@ 6.1 Datasets and Classification Models
Architecture. We use two architectures:(i) a five-layer CNN architecture with five standard convolution layers with two $3\times 3$ filters and ReLU activation function, followed by three fully connected layers; and (ii) a ResNet18. Refer to Table 5 in the Appendix for the details of these architectures.
@@ 6.2 Static Backdoor Attack
Static-backdoor Task. For this experiment, we implanted the backdoor in the first task by selecting 5% of the training samples. We inserted a $4\times 4$-pixel static trigger, with pixel values set to 0, at the bottom right corner.
@@
Results. Table 2 presents the results of the LTB attack (refer to Table 6 in the Appendix for BTB attack results), indicating the effectiveness of our latent backdoor against all CL algorithms; the blind backdoor exhibits similar attack efficacy. In most scenarios, the proposed LTB attack achieves attack success rates (ASRs) above 90%, with the lowest ASR being 82.02% for the attack on ResNet18-SplitCIFAR10 using XdG. We summarized these results in Figure 7. One can observe that when using CNN architecture, the LTB attack against regularization-based and replay-based algorithms achieves an average ASR of 90.6% and 90.7%, respectively. For the CNN architecture, the LTB attack has led to minimal accuracy drops of 5.1% for regularization-based algorithms, but a negligible classification accuracy increase of 0.8% for replay-based algorithms. For the ResNet18 architecture, the LTB attack against regularization-based and replay-based algorithms achieves an average ASR of 90.8% and 89.4%, respectively. Moreover, the attack minimally reduced the classification accuracy by 4.3% and 6.5% for regularization-based and replay-based algorithms, respectively. These results confirm the effectiveness of the LTB against continual learning using static triggers.
@@ 6.3 Dynamic Backdoor Attack
Main task. For this experiment, we use the ResNet18 model with SplitCIFAR10 and employ the LwF and DGR algorithms as representatives of regularization-based and replay-based approaches, respectively.
Dynamic trigger Task. We uniformly selected 15% of the training samples from task 1, ensuring an equal number of images with label 0 (airplane) and label 1 (car). A $5\times 5$ trigger with random pixel values was then inserted at random positions on the selected inputs. Compared to the static trigger, we slightly increased the ratio of triggered images to compensate for the dynamic nature of the trigger’s position and color.
@@ 6.4 Physical Backdoor Attack
Figure 10: LTB performance using physical triggers for LwF and DGR, with ResNet18 on SplitCIFAR10. Despite the challenges posed by physical triggers, the attack achieved a comparable ASR to the static trigger, with nearly identical ACC (less than a 1% drop).
@@ 6.5 Words Backdoor Attack
Figure 13: Under the LwF and DGR algorithm, our proposed LTB and BTB backdoor attacks demonstrate higher efficacy and persistence compared to other backdoor attacks, including BadNets [17], dynamic attack  [39], and clean-label attack. [54]. The ASR values of our BTB and LTB are higher than those of others, and the ACC values of different models are close to the baseline value (without attack).
@@ 6.7 Attacking Different Tasks
Backdoor different tasks. We apply a $4\times 4$ static trigger to 5% of the data in each dataset to backdoor the model and evaluate individual LTB attacks initiated at different tasks.
Results. Figure 15 shows the ASR and ACC, with each line representing the attack initiated at a specific task. For EWC on SplitMNIST, the attack success rates remain within the range of $0.89\%$ to $0.99\%$ across all tasks, with an average ASR drop of less than 2% per task. The ACC values consistently remain above $83\%$. For EWC on PermutedMNIST, we evaluated the attack initiated at tasks 1, 3, 5, 7, and 9. The ASR values range from $0.81\%$ to $0.97\%$, with an average ASR drop of less than 1.4% per task. The ACC across all tasks remains above $76\%$. For A-GEM on SplitMNIST, we observed consistent ASRs in the range of $0.94\%$ to $0.98\%$, with less than a 0.8% ASR drop per task, on average. For A-GEM on PermutedMNIST, the ASR values range from $0.89\%$ to $0.93\%$, leading to an average ASR drop of 0.8% per task. Across the 5 tasks for SplitMNIST, the ACC drops by only 0.8% per task, and for PermutedMNIST, which has 10 tasks, the ACC drops by just 1.2% per task.
@@ 6.8 Evading Defenses
Results. Table 3 presents the results with and without applying the defenses. For the SentiNet defense, the average ACC of the LwF algorithm across five tasks drops by 2.4%, with a negligible ASR decrease of 0.56%. Under the I-BAU defense, the average ACC drops by $4.8\%$, and the ASR decreases by only 2.2%. These results demonstrate that our LTB effectively evades both defenses.
@@ 7 Conclusion
In this paper, we demonstrated that existing backdoor attacks, e.g., clean-label and dynamic backdoors, are ineffective in continual learning settings, as changes to the model’s parameters during the learning of new data distributions diminish the backdoor’s impact. We proposed two persistent backdoor attacks: Blind Task Backdoor and Latent Task Backdoor. Our blind attack subtly alters loss computation across tasks without the adversary controlling the training process. Our latent attack embeds backdoors into the most stable components of a single task, ensuring that the malicious behavior persists even after multiple tasks. We conducted extensive evaluations using a variety of triggers–including static, dynamic, physical, and semantic triggers–across multiple continual learning algorithms. Our results demonstrate that both propose attacks consistently achieve high success rates and maintain as the model evolves. We also showed that persistent backdoor attacks effectively evade state-of-the-art detection defenses, e.g., SentiNet and I-BAU. Highlighting the needs for a new breed of backdoor detection methodology, which we plan to explore.
@@ 8 Ethical Considerations
9 Compliance with Open Science Policy
In alignment with USENIX Security’s Open Science policy, we commit to openly sharing all research artifacts associated with this work. This includes the complete codebase, raw and processed data, and all datasets used for experimentation, as well as scripts required to reproduce the results presented in the paper. Our aim is to enhance the reproducibility and transparency of our research, allowing the broader community to replicate and build upon our findings. All materials will be made available upon paper acceptance, in compliance with the artifact evaluation process.

### v3

@@ 6.1 Datasets and Classification Models
Architecture. We use two architectures:(i) a five-layer CNN architecture with five standard convolution layers with two $\times 3$ 3 filters and ReLU activation function, followed by three fully connected layers; and (ii) a ResNet18. Refer to Table 5 in the Appendix for the details of these architectures.
@@ 6.2 Static Backdoor Attack
Static-backdoor Task. For this experiment, we implanted the backdoor in the first task by selecting 5% of the training samples. We inserted a $\times 4$ 4-pixel static trigger, with pixel values set to 0, at the bottom right corner.
Results. Table 2 presents the results of the LTB attack (refer to Table 6 in the Appendix for BTB attack results), indicating the effectiveness of our latent backdoor against all CL algorithms; the blind backdoor exhibits similar attack efficacy. In most scenarios, the proposed LTB attack achieves attack success rates (ASRs) above 90%, with the lowest ASR being 82.02% for the attack on ResNet18-SplitCIFAR10 using XdG. We summarized these results in Figure 9. One can observe that when using CNN architecture, the LTB attack against regularization-based and replay-based algorithms achieves an average ASR of 90.6% and 90.7%, respectively. For the CNN architecture, the LTB attack has led to minimal accuracy drops of 4.6% for regularization-based algorithms, but a negligible classification accuracy increase of 0.6% for replay-based algorithms. For the ResNet18 architecture, the LTB attack against regularization-based and replay-based algorithms achieves an average ASR of 90.8% and 89.4%, respectively. Moreover, the attack minimally reduced the classification accuracy by 3.9% and 5.2% for regularization-based and replay-based algorithms, respectively. These results confirm the effectiveness of the LTB against continual learning using static triggers.
@@
We also analyzed the loss trajectory of ResNet18 on the validation set of SplitCIFAR10 using A-GEM, attacking Task 1 with the trigger embedded in stable and random neurons (Figure 10). The bandwidth graphs are generated from twenty experiments conducted with different random seeds. When embedding the attack in the stable neurons of Task 1, both the target and attack losses remain consistently low across all tasks, indicating attack persistence. In contrast, embedding the attack in randomly selected neurons results in a significant increase in attack loss starting from Task 2, indicating attack failure. This behavior is due to significant changes in the values of randomly selected neurons, underscoring the importance of neuron stability for attack effectiveness. We refer readers to Figure 20 in the Appendix 10 for training loss.
Figure 10: Loss trajectory comparison of attack embedding in stable versus random neurons in Task 1 reveals distinct behaviors. Starting from Task 2, the attack loss remains consistently low with stable neurons, while using random neurons leads to a sharp increase, rendering the attack ineffective.
@@ 6.3 Dynamic Backdoor Attack
Main task. We use the ResNet18 model with SplitCIFAR10 and employ the LwF and DGR algorithms for regularization-based and replay-based approaches, respectively.
Dynamic trigger Task. We uniformly selected 15% of the training samples from task 1, ensuring an equal number of images with label 0 (airplane) and label 1 (car). A $\times 5$ 5 trigger with random pixel values was then inserted at random positions on the selected inputs. Compared to the static trigger, we slightly increased the ratio of triggered images to compensate for the dynamic nature of the trigger’s position and color.
@@ 6.4 Physical Backdoor Attack
Figure 13: LTB performance using physical triggers for LwF and DGR, with ResNet18 on SplitCIFAR10. Despite the physical trigger challenges, the attack achieved a comparable ASR and ACC to the static trigger (less than 1% ACC drop).
@@ 6.5 Words Backdoor Attack
Figure 16: Under the LwF and DGR algorithm, our proposed LTB and BTB backdoor attacks demonstrate higher efficacy and persistence compared to other backdoor attacks, including BadNets [17], dynamic attack  [39], and clean-label attack [56]. The ASR values of our BTB and LTB are higher than those of others, and the ACC values of different models are close to the baseline value (without attack).
@@ 6.7 Attacking Different Tasks
Backdoor different tasks. We apply a $\times 4$ 4 static trigger to 5% of the data in each dataset to backdoor the model and evaluate individual LTB attacks initiated at different tasks.
[…]
@@ 6.8 Evading Defenses
[…]
@@ 7 Conclusion
We demonstrated that existing backdoor attacks are ineffective in CL settings, as changes to the model’s parameters during the learning of new data distributions diminish the backdoor’s impact. We propose two persistent backdoor attacks: Blind Task Backdoor and Latent Task Backdoor. The former subtly modifies loss computation across tasks without requiring adversarial control over the training process, while the latter embeds backdoors into the most stable components of a single task, ensuring persistent adversarial behavior. We conducted extensive evaluations using a variety of triggers across multiple CL algorithms. Our results demonstrate that both propose attacks consistently achieve high efficacy and maintain as the model evolves. Our attacks effectively evade existing detection defenses, e.g., SentiNet and I-BAU, highlighting the needs for more effective detection methodologies.
Acknowledgements
This research was partially funded by the US National Science Foundation under grants #2148358 and #2133407. Any opinions, findings, or conclusions expressed in this material are those of the authors and do not necessarily reflect the views of the US federal agencies.
@@ 8 Ethical Considerations
9 Open Sciences
In alignment with USENIX Security’s Open Science policy, we aim to enhance the reproducibility and transparency of our research, allowing the broader community to replicate and build upon our findings. Therefore, we share the artifacts associated with this work, including the complete codebase, raw and processed data that are not publicly available but used in our experimentation, as well as scripts required to reproduce the results presented in the paper. In compliance with the artifact evaluation process, all materials are available at: https://doi.org/10.5281/zenodo.14728872.